# Symmetry Breaking in an Extended O(2) Model

### 29 Dec 2023 v1

@@ Abstract
Motivated by attempts to quantum simulate lattice models with continuous Abelian symmetries using discrete approximations, we study an extended-O(2) model in two dimensions that differs from the ordinary O(2) model by the addition of an explicit symmetry breaking term $-h_{q}\cos(q\varphi)$. Its coupling $h_{q}$ allows to smoothly interpolate between the O(2) model ($h_{q}=0$) and a $q$-state clock model ($h_{q}\rightarrow\infty$). In the latter case, a $q$-state clock model can also be defined for non-integer values of $q$. Thus, such a limit can also be considered as an analytic continuation of an ordinary $q$-state clock model to noninteger $q$. In previous work, we established the phase diagram of the model in the infinite coupling limit ($h_{q}\rightarrow\infty$). We showed that for non-integer $q$, there is a second-order phase transition at low temperature and a crossover at high temperature. In this work, we establish the phase diagram at finite values of the coupling using Monte Carlo and tensor methods. We show that for non-integer $q$, the second-order phase transition at low temperature and crossover at high temperature persist to finite coupling. For integer $q=2,3,4$, we know there is a second-order phase transition at infinite coupling (i.e. the well-known clock models). At finite coupling, we find that the critical exponents for $q=3,4$ vary with the coupling, and for $q=4$ the transition may turn into a BKT transition at small coupling. We comment on the similarities and differences of the phase diagrams with those of quantum simulators of the Abelian-Higgs model based on ladder-shaped arrays of Rydberg atoms.
@@ I Introduction
In [7, 19], we studied the $h_{q}\rightarrow\infty$ limit of this model [^1] The action Eq. (II) differs slightly from the convention of [7], where $\beta$ was treated as a coupling attached to the first term in the action. In the present work, $\beta$ is factored out of the action, so that it plays the role of inverse temperature.. For non-integer $q$, we found a crossover at small $\beta$ and a second-order phase transition of the 2D Ising universality class at larger $\beta$. Thus we established the phase diagram for the case $h_{q}\rightarrow\infty$. The phase diagram for $h_{q}=0$ and $h_{q}=\infty$ is shown in the left panel of Fig. 1. In the present work, we study the phase diagram for finite $h_{q}$. Preliminary results were presented in [20]. We performed a scan of the parameter space and computed the specific heat. In many second-order phase transitions the specific heat diverges near the critical point. For such a transition, which is characterized by critical exponent $\alpha>0$, a heatmap of the specific heat can serve as a proxy for the phase diagram. In the right panel of Fig. 1, we insert several two-dimensional slices where each slice shows a heatmap of the specific heat at a fixed $h_{q}$. The heatmaps were computed using tensor renormalization group (TRG) methods developed for the O(2) model [21, 22, 23, 24] with size $L=1024$. The resulting picture suggests that the sharp edges and corners that one sees at integer values of $q$ in the $h_{q}=\infty$ phase diagram become smooth curves at finite $h_{q}$. However, as we will show later, this does not hold up under scrutiny.
@@
Figure 1: We consider the three-dimensional phase diagram of the Extended-O(2) model for parameter values $h_{q},\beta\in\mathbb{R}\geq 0$ and $q\in\mathbb{R}$. In the $h_{q}=0$ plane, the model reduces to the $XY$ model for all values of $q$. The $XY$ model has a disordered phase at small $\beta$, a BKT transition near $\beta_{c}=1.12$, and a critical phase at large $\beta$. We previously established the phase diagram of the $h_{q}=\infty$ plane [7]. For integer $q$, it is the well-studied clock model. For noninteger $q$, there is a crossover at small $\beta$ and a second-order phase transition of the 2D Ising universality class at large $\beta$. Establishing the phase diagram at finite-$h_{q}$ is the goal of the present work. We performed a scan of the parameter space and computed the specific heat. For phase transitions in which the specific heat diverges near the critical point, a heatmap of the specific heat can serve as a proxy for the phase diagram. In the figure on the right, we insert four two-dimensional slices where each slice shows a heatmap of the specific heat at a fixed $h_{q}$. The heatmaps were computed using TRG with $L=1024$. The resulting picture suggests that the sharp edges and corners that one sees at integer values of $q$ in the $h_{q}=\infty$ phase diagram become smooth curves at finite $h_{q}$. However, as we will show later, this does not hold up under scrutiny.
@@
Specific cases of the Extended-O(2) model have been studied before. For $q=2,3$, an early normalization group analysis [8] showed that the model at finite $h_{q}$ should be in the Ising and 3-state Potts universality class respectively. However, Ref. [9] demonstrated that the critical exponents vary with $h_{q}$, and concluded that for $q=3$, there is a BKT phase for sufficiently small $h_{q}$. The $q=4$ with finite $h_{q}$ case, sometimes referred to as the $XYh_{4}$ model, has been studied extensively due to its possible relevance to two-dimensional physical systems with 4-fold symmetry such as certain magnetic thin films [10, 11]. The renormalization group analysis [8] showed there to be a second-order phase transition with nonuniversal critical exponents which vary with $h_{q}$. The BKT phase of the O(2) model was recovered only for $h_{q}=0$. In Ref. [12], numerical MC studies of the $q=4$ case showed that the critical exponent $\nu$ does vary with $h_{q}$. This was supported further by Refs. [13, 14], and more discussion can be found in Refs. [10, 15]. However, in Ref. [16], it was found that for sufficiently small $h_{q}$, the second-order phase transition for $q=4$ is replaced by BKT transitions. This is supported by Ref. [9]. However, recent work [11, 17] suggests that the apparent BKT phase at finite $h_{q}$ is essentially a finite-size effect and might not persist to infinite volumes. For $q\geq 5$, the renormalization group analysis [8], showed there should be a BKT phase between a low-temperature ferromagnetic phase and a high-temperature paramagnetic phase. For $q=6$, this is supported by numerical MC studies on triangular lattices [14]. The $q=5$ case was shown to have two transitions for any finite $h_{q}$ [18].
@@ II The model
When $h_{q}=0$, this is the ordinary O(2) model. When $h_{q}>0$, the O(2) symmetry is broken. When $h_{q}\rightarrow\infty$ and $q$ is integer, the spin angles are forced to take the values $\varphi^{(k)}_{x}=2\pi k/q$ with $k\in\mathbb{Z}$, and hence this becomes the $q$-state clock model. We note that the action defined in Eq. (II) is also valid for non-integer $q$.
For non-integer $q$, the $2\pi$-rotational symmetry is broken so we have to carefully define the domain of the angles $\varphi_{x}$. For example, one could choose $\varphi_{x}\in[-\pi,\pi)$ or $\varphi_{x}\in[0,2\pi)$. These different choices lead to very different phase diagrams as we showed for the $h_{q}\rightarrow\infty$ case in Ref. [7]. We prefer the choice $\varphi_{x}\in[0,2\pi)$, simply because the phase diagram gives a more consistent periodic picture as $q$ is increased [^2] See the left panel of Figure 26 in Ref. [7] as opposed to the right panel in the same figure..
@@ IV.1 Integer $q$
We summarize now the results for integer $q$. Recall that for $h_{q}=\infty$, this model becomes the ordinary $q$-state clock model, which is known to have a second-order transition for $q=2,3,4$, and a pair of BKT transitions for $q\geq 5$. In this work, we studied what happens at finite $h_{q}$. For $q\geq 5$, there seems to be a pair of BKT transitions for all values of $h_{q}>0$. For $q=2,3,4$, there is a second-order transition for large and intermediate values of $h_{q}$. For $q=3,4$, the critical exponents, and hence the universality class, vary with $h_{q}$, and for $q=4$ this transition appears to turn into a BKT transition at small $h_{q}$. There are very few results in the literature even for integer $q$ (we found no literature on noninteger $q$) that give specific numbers such as critical exponents or critical points. Of the few results that we found, our findings seem to agree with, for example, [14, 16, 11] but disagree with [9]. The reason for this is not clear to us.
@@ IV.2 Noninteger $q$
At noninteger $q$, the explicitly broken $\mathbb{Z}_{q}$ symmetry causes difficulties for the MCMC approach. In fact, we were unable to get reliable results on lattices larger than $32\times 32$ due to the large autocorrelations that result. This makes finite-size scaling infeasible with the MCMC approach. Instead, we use TRG to investigate the non-integer $q$ regime. We validate the TRG approach by comparison with MCMC on smaller lattices. See Appendix D.
With TRG, we started with a low precision scan of the parameter space for $q\in[1,6]$, $\beta\in[0,2]$, $h_{q}=0.1,1.0$, and $L=4,8,\ldots,1024$. For example, we show in Fig. 6 heatmaps of the specific heat (top row) and entanglement entropy (bottom row) for $h_{q}=1$ (left column) and $h_{q}=0.1$ (right column). Each pixel is a data point obtained from TRG performed with $L=1024$ and bond dimension 40. The color in each heatmap ranges from dark to light and this corresponds to a value ranging from 0 to 2.5. The cutoff choice of 2.5 (which truncates some of the specific heat values) was made to increase the contrast in the heatmaps. For $h_{q}=1$, the heatmap of the specific heat shows smooth lobes suggesting smooth lines in the phase diagram even as one crosses integer values of $q$. When $h_{q}$ is reduced to 0.1, the lines at large $\beta$ fade away leaving a thick horizontal line that connects to the BKT transition at $h_{q}=0$. However, heatmaps of the entanglement entropy show that there remain obvious discontinuities as one crosses integer values of $q$. This suggests a phase diagram at finite $h_{q}$ that is similar to the phase diagram at infinite $h_{q}$.
At noninteger values of $q$, the specific heat generally shows two peaks. In the $h_{q}=\infty$ limit, it was found [7] that the first peak is only a crossover, but the second peak is associated with a second-order phase transition of the 2D Ising universality class. Here, we investigate the finite $h_{q}$ regime.
@@ IV.3 Phase Diagram
In the limit $h_{q}\rightarrow\infty$, for $q=2,3,4$, there is a single second-order phase transition of the $q$-state clock model universality class, and for integer $q\geq 5$, there are two BKT phase transitions. When $h_{q}$ is dialed toward zero, the two BKT transitions for $q\geq 5$ appear to persist to all $h_{q}>0$. The small-$\beta$ BKT transitions merge with the BKT transition of the $XY$ model at $h_{q}=0$. The large-$\beta$ BKT transitions, which vanish only for $h_{q}=0$, converge to different but finite values of $\beta$ in the limit $h_{q}\rightarrow 0$ [18]. For $q=3,4$, the critical exponents of the second-order phase transitions vary with $h_{q}$. We see evidence that the transition for $q=4$ turns into a BKT phase transition at small $h_{q}>0$.

### 26 Feb 2024 v2

@@ Abstract
Motivated by attempts to quantum simulate lattice models with continuous Abelian symmetries using discrete approximations, we study an extended-O(2) model in two dimensions that differs from the ordinary O(2) model by the addition of an explicit symmetry breaking term $-h_{q}\cos(q\varphi)$. Its coupling $h_{q}$ allows to smoothly interpolate between the O(2) model ($h_{q}=0$) and a $q$-state clock model ($h_{q}\rightarrow\infty$). In the latter case, a $q$-state clock model can also be defined for noninteger values of $q$. Thus, such a limit can also be considered as an analytic continuation of an ordinary $q$-state clock model to noninteger $q$. In previous work, we established the phase diagram for noninteger $q$ in the infinite coupling limit ($h_{q}\rightarrow\infty$). We showed that there is a second-order phase transition at low temperature and a crossover at high temperature. In this work, we seek to establish the phase diagram at finite values of the coupling using Monte Carlo and tensor methods. We show that for noninteger $q$, the second-order phase transition at low temperature and crossover at high temperature persist to finite coupling. For integer $q=2,3,4$, we know there is a second-order phase transition at infinite coupling (i.e. the well-known clock models). At finite coupling, we find that the critical exponents for $q=3,4$ vary with the coupling, and for $q=4$ the transition may turn into a BKT transition at small coupling. We comment on the similarities and differences of the phase diagrams with those of quantum simulators of the Abelian-Higgs model based on ladder-shaped arrays of Rydberg atoms.
@@ I Introduction
In [7, 19], we studied the $h_{q}\rightarrow\infty$ limit of this model [^1] The action Eq. (II) differs slightly from the convention of [7], where $\beta$ was treated as a coupling attached to the first term in the action. In the present work, $\beta$ is factored out of the action, so that it plays the role of inverse temperature.. For integer values of $q$, this is the well-studied clock model. For noninteger $q$, we found a crossover at small $\beta$ and a second-order phase transition of the 2D Ising universality class at larger $\beta$. Thus we extended the phase diagram for the case $h_{q}\rightarrow\infty$ to noninteger values of $q$. The phase diagram for $h_{q}=0$ and $h_{q}=\infty$ is shown in the left panel of Fig. 1. In the present work, we study the phase diagram for finite $h_{q}$. Preliminary results were presented in [20]. We performed a scan of the parameter space and computed the specific heat. In many second-order phase transitions the specific heat diverges near the critical point. For such a transition, which is characterized by critical exponent $\alpha>0$, a heatmap of the specific heat can serve as a proxy for the phase diagram. In the right panel of Fig. 1, we insert several two-dimensional slices where each slice shows a heatmap of the specific heat at a fixed $h_{q}$. The heatmaps were computed using tensor renormalization group (TRG) methods developed for the O(2) model [21, 22, 23, 24] with size $L=1024$. The resulting picture suggests that the sharp edges and corners that one sees at integer values of $q$ in the $h_{q}=\infty$ phase diagram become smooth curves at finite $h_{q}$. However, as we will show later, this does not hold up under scrutiny.
@@
Figure 1: (Left) We consider the three-dimensional phase diagram of the Extended-O(2) model for parameter values $h_{q},\beta\in\mathbb{R}\geq 0$, and $q\in\mathbb{R}$. In the $h_{q}=0$ plane, the model reduces to the $XY$ model for all values of $q$. The $XY$ model has a BKT transition near $\beta_{c}=1.12$ and a critical phase for $\beta>\beta_{c}$. In the $h_{q}=\infty$ plane, the phase diagram is more interesting. For integer $q$, the model reduces to the well-studied clock model, which has a second-order phase transition for $q=2,3,4$ and two BKT transitions for integer $q\geq 5$. For noninteger $q$, we previously established the phase diagram for $h_{q}\rightarrow\infty$ [7]. We found crossovers (dashed lines) at small $\beta$ and second-order phase transitions (solid black lines) of the 2D Ising universality class at large $\beta$. Establishing the phase diagram at finite-$h_{q}$ is the goal of the present work. (Right) We insert several two-dimensional slices where each slice shows a heatmap, computed using TRG with $L=1024$, of the specific heat at a fixed $h_{q}$. As described in the main text, such heatmaps can serve as a proxy for the phase diagram. Here they suggest that the sharp edges and corners seen in the $h_{q}=\infty$ phase diagram become smooth curves at finite $h_{q}$. However, as we will show later, this does not hold up under scrutiny. A finite-size scaling analysis will show that for noninteger $q$, the phase diagram looks qualitatively the same for all $h_{q}>0$.
@@
Specific cases of the Extended-O(2) model have been studied before. For $q=2,3$, an early normalization group analysis [8] showed that the model at finite $h_{q}$ should be in the Ising and 3-state Potts universality class respectively. However, Ref. [9] demonstrated that the critical exponents vary with $h_{q}$, and concluded that for $q=3$, there is a BKT phase for sufficiently small $h_{q}$. The $q=4$ with finite $h_{q}$ case, sometimes referred to as the $XYh_{4}$ model, has been studied extensively due to its possible relevance to two-dimensional physical systems with 4-fold symmetry such as certain magnetic thin films [10, 11]. The renormalization group analysis [8] showed there to be a second-order phase transition with nonuniversal critical exponents which vary with $h_{q}$. The BKT phase of the O(2) model was recovered only for $h_{q}=0$. This is supported by more recent work using a nonperturbative renormalization group approach wherein the model was studied with a continuously varying spatial dimensionality $2\leq D\leq 3$ [30]. In Ref. [12], numerical MC studies of the $q=4$ case showed that the critical exponent $\nu$ does vary with $h_{q}$. This was supported further by Refs. [13, 14], and more discussion can be found in Refs. [10, 15]. However, in Ref. [16], it was found that for sufficiently small $h_{q}$, the second-order phase transition for $q=4$ is replaced by BKT transitions. This is supported by Ref. [9]. However, recent work [11, 17] suggests that the apparent BKT phase at finite $h_{q}$ is essentially a finite-size effect and might not persist to infinite volumes. For $q\geq 5$, the renormalization group analysis [8], showed there should be a BKT phase between a low-temperature ferromagnetic phase and a high-temperature paramagnetic phase. For $q=6$, this is supported by numerical MC studies on triangular lattices [14]. The $q=5$ case was shown to have two transitions for any finite $h_{q}$ [18].
@@
The goal of the present work is to establish the phase diagram of the Extended-O(2) model at finite $h_{q}$ for both integer and noninteger values of $q$. This is motivated by several things. To quantum simulate Abelian lattice models, it may be necessary to approximate U(1) with, e.g., discrete $\mathbb{Z}_{q}$ approximations. The Extended-O(2) model with integer $q$ allows to smoothly break the O(2) symmetry of the O(2) model down to the $\mathbb{Z}_{q}$ symmetry of a clock model. Studying the effect of this symmetry-breaking is crucial to understanding the utility of various $\mathbb{Z}_{q}$ approximations of $U(1)$. Furthermore, by allowing $q$ to be noninteger, we can explore the effect of breaking the $\mathbb{Z}_{q}$ symmetry itself, which may be useful for understanding the resilience of a real-life $\mathbb{Z}_{q}$ approximation to perturbations. Finally, early results from our study of this model suggested a phase diagram reminiscent of one from Rydberg atom chains where exotic floating phases are found. By studying the phase diagram of the Extended-O(2) model at finite $h_{q}$, we will determine if similar exotic phases can be found in this classical spin system.
@@ II The model
When $h_{q}=0$, this is the ordinary O(2) model. When $h_{q}>0$, the O(2) symmetry is broken. When $h_{q}\rightarrow\infty$ and $q$ is integer, the spin angles are forced to take the values $\varphi^{(k)}_{x}=2\pi k/q$ with $k\in\mathbb{Z}$, and hence this becomes the $q$-state clock model. We note that the action defined in Eq. (II) is also valid for noninteger $q$.
For noninteger $q$, the $2\pi$-rotational symmetry is broken so we have to carefully define the domain of the angles $\varphi_{x}$. For example, one could choose $\varphi_{x}\in[-\pi,\pi)$ or $\varphi_{x}\in[0,2\pi)$. These different choices lead to very different phase diagrams as we showed for the $h_{q}\rightarrow\infty$ case in Ref. [7]. We prefer the choice $\varphi_{x}\in[0,2\pi)$, simply because the phase diagram gives a more consistent periodic picture as $q$ is increased [^2] See the left panel of Figure 26 in Ref. [7] as opposed to the right panel in the same figure..
@@ IV.1 Integer $q$
We summarize now our results for integer $q$ and compare them with prior literature. In the infinite-coupling limit $h_{q}\rightarrow\infty$, this model becomes the ordinary $q$-state clock model. For integer $q$, the clock model has been well-studied and is known to have a second-order transition for $q=2,3,4$ and a pair of BKT transitions for $q\geq 5$. In our previous work [7], we studied this limit of the model, and our results for integer $q$ were consistent with the prior literature. In the present work, we study the model at finite $h_{q}$. Specific cases appear in the literature, but nobody had yet performed a study of the critical exponents of the model across a large parameter range. For $q=2$, our results are mostly consistent with an Ising phase transition for the entire range of $h_{q}$ that we studied. This is consistent with an early renormalization group analysis [8]. For $q=3$, Ref. [8] predicted the 3-state Potts universality class for all $h_{q}>0$. Our numerical results agree with this for large $h_{q}$ but not for small $h_{q}$, where we find a second-order phase transition with critical exponents different from the 3-state Potts model. Our results for $q=3$ and small $h_{q}$ are more consistent with a numerical study [9], which showed that the critical exponents vary with $h_{q}$. This study also concluded that there is a BKT phase for values of $h_{q}$ smaller than what we studied here. For $q=4$ with large and intermediate values of the coupling, we find a second-order phase transition with critical exponents that vary with $h_{q}$. This is consistent with prior literature [8, 30, 12, 13, 14, 10, 15]. The literature agrees on what happens at large and intermediate $h_{q}$, but disagrees on what happens at small $h_{q}$. The early renormalization group analysis [8] showed that the BKT transition is recovered only for $h_{q}=0$. Numerical studies [16, 9] found evidence for BKT transitions at small $h_{q}$, however, more recent work [11, 17] suggests that the apparent BKT phase at finite $h_{q}$ is essentially a finite-size effect and might not persist to infinite volumes. Our own results for $q=4$ and small $h_{q}=0.1$ are consistent with BKT, however, our lattices are not large enough to exclude the possibility that these are finite-size effects. Only a few studies by other groups give quantitative results, such as critical points and exponents, which we can compare against. For $q=4$ and $h_{q}=1.0$, our critical point and critical exponents seem consistent with those in Refs. [14, 11]. For $q=4$ and $h_{q}=0.1$, our results for the critical point and critical exponent $\beta$ agree with those in Ref. [16]. For $h_{q}=0.1$ our quantitative results for $\nu$ and $\gamma$ disagree with those in Ref. [9], although we agree with the conclusion that the behavior is BKT-like. For integer $q\geq 5$, our results suggest a pair of BKT transitions for all values of $h_{q}>0$. This is consistent with prior literature including renormalization group analysis [8] and numerical studies of specific cases [14, 18].
@@ IV.2 Noninteger $q$
At noninteger $q$, the explicitly broken $\mathbb{Z}_{q}$ symmetry causes difficulties for the MCMC approach. In fact, we were unable to get reliable results on lattices larger than $32\times 32$ due to the large autocorrelations that result. This makes finite-size scaling infeasible with the MCMC approach. Instead, we use TRG to investigate the noninteger $q$ regime. We validate the TRG approach by comparison with MCMC on smaller lattices. See Appendix D.
With TRG, we started with a low precision scan of the parameter space for $q\in[1,6]$, $\beta\in[0,2]$, $h_{q}=0.1,1.0$, and $L=4,8,\ldots,1024$. For example, we show in Fig. 6 heatmaps of the specific heat (top row) and entanglement entropy (bottom row) for $h_{q}=1$ (left column) and $h_{q}=0.1$ (right column). Each pixel is a data point obtained from TRG performed with $L=1024$ and bond dimension 40. The color in each heatmap ranges from dark to light and this corresponds to a value ranging from 0 to 2.5. The cutoff choice of 2.5 (which truncates some of the specific heat values) was made to increase the contrast in the heatmaps. For $h_{q}=1$, the heatmap of the specific heat shows smooth lobes even as one crosses integer values of $q$. When $h_{q}$ is reduced to 0.1, the lines at large $\beta$ fade away leaving a thick horizontal line that connects to the BKT transition at $h_{q}=0$. Heatmaps of the entanglement entropy (bottom row of Fig. 6) show that there remain obvious discontinuities as one crosses integer values of $q$.
The heatmaps of the specific heat show that the heights and positions of the peaks in the specific heat vary smoothly as $q$ is varied. The resulting smooth lobes suggest smooth lines in the phase diagram even as one crosses integer values of $q$. Such a phase diagram is reminiscent of that of Rydberg atom chains as shown in Fig. 2 and is a motivation for carefully checking for deeper similarities between these two models. On the other hand, heatmaps of the entanglement entropy show that there remain obvious discontinuities as one crosses integer values of $q$. This suggests a phase diagram at finite $h_{q}$ that is similar to the phase diagram at infinite $h_{q}$. A rigorous finite-size scaling analysis is needed to establish the true phase diagram at noninteger $q$, and that is what we present in the remainder of this section.
At noninteger values of $q$, the specific heat generally shows two peaks. In the $h_{q}=\infty$ limit, it was found [7] that the small-$\beta$ peak is only a crossover, but the large-$\beta$ peak is associated with a second-order phase transition of the 2D Ising universality class. Here, we investigate the finite $h_{q}$ regime.
@@
In summary, for all noninteger $q$ and all $h_{q}>0$, we find a crossover at small-$\beta$ and a phase transition of the Ising universality class at large $\beta$. As $q$ approaches an integer from below, the Ising transition line smoothly connects to the critical point of the model with integer $q$. However, for $q$ a little larger than the integer, the Ising transition occurs at much larger $\beta$. The apparent similarity with the phase transition of a Rydberg atom chain therefore does not hold up under scrutiny. In fact, the true phase diagram for noninteger $q$ looks qualitatively the same for all $h_{q}>0$ with sharp discontinuities at integer values of $q$.
@@ IV.3 Phase Diagram
In the limit $h_{q}\rightarrow\infty$, for $q=2,3,4$, there is a single second-order phase transition of the $q$-state clock model universality class, and for integer $q\geq 5$, there are two BKT phase transitions. When $h_{q}$ is dialed toward zero, the two BKT transitions for $q\geq 5$ appear to persist to all $h_{q}>0$. The small-$\beta$ BKT transitions merge with the BKT transition of the $XY$ model at $h_{q}=0$. The large-$\beta$ BKT transitions, which vanish only for $h_{q}=0$, converge to different but finite values of $\beta$ in the limit $h_{q}\rightarrow 0$ [18]. For $q=3,4$, the critical exponents of the second-order phase transitions vary with $h_{q}$. We see evidence that the transition for $q=4$ turns into a BKT phase transition at small $h_{q}>0$. There is some controversy about this point with some authors claiming there is a BKT transition at small $h_{q}$ [16, 9] and others arguing that the BKT transition only occurs for $h_{q}=0$ [8, 11, 17]. In the present work, our lattices are not large enough to conclusively answer this question.